IR PHOTODEPLETION SPECTROSCOPY OF Li(NH_{3})_{n} CLUSTERS FOR n = 4-7

\underline{TOM E. SALTER}, COREY J. EVANS, and ANDREW M. ELLIS, DEPARTMENT OF CHEM-ISTRY, UNIVERSITY OF LEICESTER, UNIVERSITY ROAD, LEICESTER, LE1 7RH, UK.; VICTOR A. MIKHAILOV, UNIVERSITY OF BIRMINGHAM, EDGBASTON, BIRMINGHAM B15 2TT, UK..

Mass-selected infrared spectra of the uncharged metal-solvent clusters Li(NH_{3})_{n} in the N-H stretching region have been recorded using IR photodepletion spectroscopy. The bands observed in the IR spectra for Li(NH_{3})_{4} have been assigned to N-H stretching vibrations from solvent molecules located in the first solvent shell. The appearance of higher frequency N-H stretching bands for n ≥ 5 is attributed to one or more ammonias in a second solvent shell. These data provide strong support for previous suggestions based on gas-phase photoionization measurements that the first solvation shell for Li(NH_{3})_{n} is complete at n = 4. They are also consistent with neutron diffraction studies of concentrated lithium-liquid ammonia solutions, where quasi-tetrahedral Li(NH_{3})_{4} is found to be the basic structural motif. Our spectra will be presented alongside computational predictions of the cluster spectra.